40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu May 8 17:00:56 2014, rana, Q, Summary, LSC, REFL_DC handoff didn't work last night 
    Reply  Fri May 9 04:51:13 2014, ericq, Summary, LSC, REFL_DC handoff didn't work last night 
Message ID: 9936     Entry time: Fri May 9 04:51:13 2014     In reply to: 9932
Author: ericq 
Type: Summary 
Category: LSC 
Subject: REFL_DC handoff didn't work last night 

Quote:

Last night after checking cabling and turning on ISS, we tried several times to handoff to REFL_DC but it didn't work at all.

Some issues:

  1. The ISS was injecting a lot of very low frequency power fluctuations because of bad AC coupling.
  2. The SR560 @ LSC rack was saturating a lot with the x10 gain that Jenne and Rana put in; we turned it back to G = 1.
  3. The ISS was also saturating a lot. We turned it off around 4 AM, but still no success.
  4. The ALS sequence for finding the Red Resonance takes too long (~2 minutes), so we're trying a faster scheme tonight.

 Still no success tonight

  • We took CARM OLTFs at various CARM offsets and could clearly see the peak in the optical TF (in once case ~2.5kHz), which gave us an indication of our offset (~200pm)
  • REFLDC effectively sees the same plant TF as the transmission signals plus a zero at ~110 Hz, at all offsets under 1nm, from my simulations; this pushes up the optical resonance and causes a loop instability when we try to handoff. 
  • We need to make the CARM OLTF steeper to suppress this instability, but also to make a good crossover with the AO path, which otherwise has too similar of a slope around the UGF, as we saw with our one arm test. 
  • We're thinking of trying to turn the AO path on with REFLDC while keeping the arms on SQRTINV signals. This may be tricky, but if we can get the loop bandwidth above the optical peak, it'll be a lot easier to deal with, and transfer digital control to REFLDC as well. 
ELOG V3.1.3-